OpenStudy (anonymous):

MEDAAAAL! Kim paid 3/7 of a bill, Sherri paid 35% of the bill, and Judy paid the rest. Which of the following statements is true? Kim paid the most. Sherri paid the most. Judy paid the most. They all paid the same amount.

OpenStudy (anonymous):

@waterineyes help?

OpenStudy (anonymous):

@robtobey @radar help?

OpenStudy (anonymous):

@skullpatrol heeelp?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Code of Conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (skullpatrol):

You need to convert the three given numerals into values that you can easily compare.

OpenStudy (anonymous):

first, you need to figure out how much 3/7 of the bill is, in percent. 3/7 *100 = ???

OpenStudy (anonymous):

that is the percent than Kim paid. you know how much sherri paid. Judy: she paid 100% - (what kim paid + what sherri paid). then just see which statement is true...

OpenStudy (anonymous):

A. Kim paid the most. 3/7 = .429 35% = .350 Left Over = .22, or 22%

OpenStudy (anonymous):

the answer is A.

OpenStudy (anonymous):

they paid the same

OpenStudy (anonymous):

Kim paid about 45% Sherri paid 35%. 45+35=78 100%-78%=22% So Kim paid the most. A is the answer